Construct a table of values of the following functions using the interval of -5 to 5



j(t)=1/t²-2t+1

1
Expert's answer
2021-11-08T06:14:08-0500
"j(t)=\dfrac{1}{t^2-2t+1}=\dfrac{1}{(t-1)^2}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 t & j(t)\\ \hline
 -5 & 1/36 \\
 \hdashline
 -9/2 & 4/121\\
 \hdashline
 -4 & 1/25 \\
 \hdashline
 -7/2 & 4/81\\
 \hdashline
 -3 & 1/16 \\
 \hdashline
 -5/2 & 4/49\\
 \hdashline
 -2 & 1/9 \\
 \hdashline
 -3/2 & 4/25\\
 \hdashline
 -1 & 1/4\\
 \hdashline
 -1/2 & 4/9\\
 \hdashline
 0 & 1\\
 \hdashline
 1/2 & 4\\
 \hdashline
 1 & undefined\\
 \hdashline
 3/2 & 4\\
 \hdashline
 2 & 1\\
 \hdashline
 5/2 & 4/9\\
 \hdashline
 3 & 1/4\\
 \hdashline
 7/2 & 4/25\\
 \hdashline
 4 & 1/9\\
 \hdashline
 9/2 & 4/49\\
 \hdashline
 5 & 1/16\\
\end{array}"